Need help verifying answers and help answering question D answer choices are in bold.

Handbrain Inc. is considering a change to activity-based product costing. The company produces two products, cell phones and tablet PCs, in a single production department. The production department is estimated to require 2,000 direct labor hours. The total indirect labor is budgeted to be $200,000.

Time records from indirect labor employees revealed that they spent 30% of their time setting up production runs and 70% of their time supporting actual production.

The following information about cell phones and tablet PCs was determined from the corporate records:

Number of
Setups
Direct Labor
Hours
Units
Cell phones 1,200 1,000 80,000
Tablet PCs 2,800 1,000 80,000
Total 4,000 2,000 160,000

If required, round all per unit answers to the nearest cent.

a. Determine the indirect labor cost per unit allocated to cell phones and tablet PCs under a single plantwide factory overhead rate system using the direct labor hours as the allocation base.

Cell phones $ 1.25 per unit
Tablet PCs $ 1.25 per unit

b. Determine the budgeted activity costs and activity rates for the indirect labor under activity-based costing. Assume two activities—one for setup and the other for production support.

Budgeted Activity Cost Activity Rate
Setup $60,000 $15 per setup
Production support $140,00 $70 per direct labor hour

c. Determine the activity cost per unit for indirect labor allocated to each product under activity-based costing.

Cell phones $1.1 per unit
Tablet PCs $1.4 per unit

d. Why are the per-unit allocated costs in (a) different from the per-unit activity cost assigned to the products in (c)? Bold are answer choices.

The per-unit indirect labor costs in (a) are distorted because Set up activity/ production is consumed by the products in a different ratio from the direct labor. The single rate system/ activity based costing method properly allocates the setup-related activity so that the cell phones/ tablet pc's receive a larger portion of the setup activity cost. The single rate system/ activity based allocates overhead only on the basis of direct labor hours. Since the direct labor hours are/are not equal for each product, the allocated indirect labor will also not be/ be equal.
Answer #1

Firstly Let us understand how the Activity based costing is different from Traditional Costing system (Single rate system)

In Traditional Costing System, Overheads (Indirect Costs) are allocated, apportioned and lastly absorbed in the cost units may be on the basis of direct material or direct labor absorption, but actually one of the product from the product list utilises may be only low proportion of indirect cost, but by allocating the Overhead on direct material or direct labor basis will lead to get the more share of Indirect cost than actually it utilised, So there is an imbalance in the allocation of Overhead, To overcome this drawback Activity based costing is introduced, in which Overheads are allocated on the basis of benefits received from indirect activities.

The figures given in the Question are:-

Products No. of Setups Direct Labor Hours Units
Cell Phones 1200 1000 80000
Tablet PCs 2800 1000 80000
Total 4000 2000 160000

Total Budgeted Indirect Labor Cost = $200,000

Utilisation of Indirect Labor Hours(As revealed By Indirect employees):-

1) 30% for Setting up Production run

2) 70% for Actual production

a) In the first question the requirement was to determine indirect labor cost per unit of Cell phones and Tablet pcs as direct labor hours as the base for allocation under single facory overhead system

So from the above given table we can understand that:-

Total direct labor Hours = 2000 Hours (1000 Hours for Cell phone and 1000 Hours for Tablet Pcs)

Total Indirect Labor Cost (Given) = $200,000

So here the Direct labor and the No.of units of both cell phone and Table Pc are in Same proportion, so the indirect cost should be same ie:-

Cell Phone = $200,000*1000 Hours/2000 Hours =$100,000 for Total Cell phones (80000 units)

so for one unit of cell phone = $100000/80000 units = $1.25 per unit

Table PC = $200,000*1000 Hours/2000 Hours =$100,000 for Total Tablet PCs (80000 units)

so for one unit of Tablet PC = $100000/80000 units = $1.25 per unit

b) Here we want to find the budget activity cost and acitivity rate for indirect labor under Activity based costing

There are two activities; 1) Setup 2) Production support

we can understand from the question that the indirect labor time is utilising 30% for setup and 70% for production as given in the question, So

Total Indirect labor cost (Budgeted) = $200,000

Setup activity:

Budget activity cost for setup (on the basis time used in this activity) = $200,000*30% = $60000

Acivity rate per setup under indirect labor = Budget activity cost for setup/Total No.of Setup

ie, $60000/4000 setups = $15 per setup

Production support activity:

Budget activity cost for production support (on the basis time used in this activity) = $200,0000*70% = $140,000

Acivity rate per Production support under indirect labor = Budget activity cost for production support/Total Direct labor Hours

ie, $140,000/2000 Hours = $70 per direct labor hour

c) Determination of activity per unit for indirect labor allocated to each product under Activity based Costing

1) Cell Phone

Total No. of units = 80000

Total Indirect cost = $200,000

Setup cost/unit = Total setup cost* No. of setup for cell phone/Total No. of setups

ie, $60000*(1200/4000)/80000 units = $.225

Production cost = $140000*(1000 Hours/2000 Hours)/80000 units = $.875

So Activity Per unit cost of cell phone = $1.1 ($.225+$.875)

2) Tablet PC

Total No. of units = 80000

Total Indirect cost = $200,000

Setup cost/unit = Total setup cost* No. of setup for Tablet PC/Total No. of setups

ie, $60000*(2800/4000)/80000 units = $.525

  Production cost = $140000*(1000 Hours/2000 Hours)/80000 units = $.875

  So Activity Per unit cost of Tablet PC = $1.4 ($.525+$.875)

d) The per unit indirect labor costs in (a) are distorted because set up activity is consumed by the products in a different ratio from the direct labor. The activity based costing method properly allocated the setup-related activity so that the tablet pc's receive a larger portion of the setup activity cost. The single rate system allocates overhead only on the basis of direct labor hours. Since the direct labor hours are equal for each product, the allocated indirect labor will also be equal